OpenStudy (anonymous):

how to solve the equation. 3x+12x=56

OpenStudy (anonymous):

15x=56 15x/15x=56/15 x=3.733333333

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

was it correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you have anymore questions?

OpenStudy (anonymous):

how can i learn the basics steps to do an equation im having a hard time

OpenStudy (anonymous):

with the one I just did? When ever you have x's you can add them together. If they are on opposite sides you want to subtract or add to get the x's on both on the same side opposite the other number. Once you do that you divide the x number by both sides and that is what gives you x.

OpenStudy (anonymous):

so what if it says 5x+15y=36

OpenStudy (anonymous):

15x=56 15x/15x=56/15 x=3.733333333

OpenStudy (anonymous):

I think I gave the correct answer! :D